# Revealing Modeshape Complexity and Sensitivity in Torsional Vibrations of Microcantilevers with Overhang- and T-shaped Geometries

### v2

@@ Abstract
The torsional vibration of atomic force microscope (AFM) cantilevers is critical for high-sensitivity measurements, yet existing models for width-varying cantilevers often rely on approximations that lead to significant discrepancies with experimental data. Unlike prior studies, this work introduces a refined analytical framework to precisely compute resonance frequencies and mode shapes—including higher-order modes—for overhang- and T-shaped microcantilevers, validated through targeted experimental comparisons. By systematically analyzing the effects of overhang length, we reveal previously unreported multi-maxima mode shapes and demonstrate how geometric tuning can controllably shift resonant frequencies. Furthermore, we establish a quantitative relationship between modal sensitivity and cantilever-surface coupling strength, providing actionable design principles for optimizing AFM cantilever performance. Our results not only reconcile theoretical predictions with experimental observations but also offer practical guidelines for tailoring cantilever geometry to achieve specific frequency responses in applications such as nanomechanical imaging and surface property mapping. This work advances the design of next-generation AFM probes by bridging the gap between analytical models and real-world operational demands.
@@ I Introduction
Atomic force microscopy (AFM) has emerged as a powerful tool for high-resolution surface characterization and single-molecule force spectroscopy, offering exceptional sensitivity and versatility across diverse environments (e.g., air, liquid, and vacuum) NeumanNatMeth08 ; Sajjadi2017 ; Karimpour_micron_2017_vibration . By monitoring cantilever deflections or shifts in resonant frequency, AFM enables the quantification of structural, thermal, and mechanical properties at the nanoscale Beytollah_JVC_2021_application ; DatJJAP23 . Recent advances in cantilever dynamics have further expanded its applications, from biomolecular interaction mapping to nanomechanical property imaging 15VyAPEX .
@@
Figure 1: A cantilever beam structure including an overhanging part of length $l_{0}$ and width $w_{0}$ is clamped at $x=0$ to the base (black region). For $w_{0}<w$, we have a T-shaped cantilever. The torsional modes are examined in the interaction with a sample via the effective interaction stiffness $\kappa_{l}$ (inset).
@@
In this work, we present a comprehensive analytical and numerical investigation of how geometric modifications in overhang- and T-shaped cantilevers alter their torsional mode shapes, resonant frequencies, and—critically—their measurement sensitivity. Unlike prior studies focusing solely on frequency tuning, we systematically quantify the influence of overhang and T-section dimensions on both dynamic response and modal sensitivity, bridging a key gap in AFM cantilever design.
@@ II Material and Methods
The cantilever is assumed to be fabricated from silicon nitride 19MohammadiUltra , consistent with conventional AFM experiments, with a length $L$ = 200–500 $\mu$m, width $w$ = 35 $\mu$m, and thickness $t$ = 1.5 $\mu$m. Building on the analytical framework developed for flexural vibrations in nonuniform cantilevers DatMRT22 ; DatCiP20 , we extend this approach to model torsional vibrations in overhang- and T-shaped structures [see Fig. 1]. Our analysis systematically investigates how variations in the overhang and T-section dimensions influence the resonant frequencies and mode shapes of the cantilevers. Therefore, a detailed analysis on the multi-mode behavior of the cantilever is required. We found that: (i) Multi-mode torsional behavior: We observe that higher-order torsional modes play a significant role in the vibration dynamics, necessitating a detailed analysis of their contributions. And (ii) Geometric sensitivity: The dimensions of the overhang and T-parts critically affect both the frequency spectrum and mode shape distribution, highlighting the importance of precise geometric control for optimal performance.
@@
where $\phi(x,t)$ is the deflection angle at position $x$ and time $t$. $G$ is the shear modulus and $\rho$ is the density of the beam. $J(x)$ and $I_{p}(x)$ are geometric functions of the beam cross section and the polar moment of inertia, respectively. Here, the cross section of the beam is a rectangle shape, hence, $J(x)=w(x)t^{3}/3$ and $I_{p}(x)=w^{3}(x)t/12$. It is shown that the width of the beam is $x$-dependent. Hence, the general solution of Eq. (1) is $\phi(x,t)=\phi(x)e^{i\omega t}$. Input it back to Eq. (1), one obtains the equations for the mode shape ($x$-dependent) and for the frequency. The mode shape equation reads,
@@
For the current cross-section of the beam, the thickness of the overhang part and the outer cantilever part are assumed to be the same while the width is steplike with $x$,
$w(x)=\begin{cases}w_{0},&\text{if $0<x\leq l_{0}$},\\ w,&\text{if $l_{0}<x\leq L$}.\end{cases}$ (3)
Based on Eq. (3), Eq. (2) is divided into two equations. The first equation describes the overhang part,
$\phi_{0}^{(2)}(x)+\gamma^{2}_{0}\phi_{0}(x)=0,$ (4)
and the second equation is for the cantilever part,
$\phi_{c}^{(2)}(x)+\gamma^{2}_{c}\phi_{c}(x)=0.$ (5)
Here, $\gamma_{0,c}=\omega\sqrt{\frac{\rho I_{p,0,c}}{GJ_{0,c}}}$ is the characteristic frequency. Now, the frequency ratio is
$\frac{\gamma_{c}}{\gamma_{0}}=\frac{w}{w_{0}}=\frac{1}{\kappa},$ (6)
or, $\gamma=\gamma_{c}=\frac{1}{\kappa}\gamma_{0}$ could be used for brevity. The solutions Eqs. (4) and (5) could be written as follows,
@@
The boundary conditions are
$\phi_{0}(0)=\left.\frac{d\phi_{c}(x)}{dx}\right|_{x=L}=0.$ (9)
The continuous conditions written at $l_{0}$ are
$\phi_{0}(l_{0})=\phi_{c}(l_{0}),$ (10)
and
$GJ_{0}\left.\frac{d\phi_{0}(x)}{dx}\right|_{x=l_{0}}=GJ_{c}\left.\frac{d\phi_{c}(x)}{dx}\right|_{x=l_{0}}.$ (11)
From these conditions, a matrix equation has been obtained as follows,
$K\cdot X=0,$ (12)
where $K$ is written as
$K=\left[\begin{matrix}0&1&0&0\\ 0&0&\cos\gamma&-\sin\gamma\\ \sin\kappa\eta\gamma&\cos\kappa\eta\gamma&-\sin\eta\gamma&-\cos\eta\gamma\\ \kappa^{2}\cos\kappa\eta\gamma&-\kappa^{2}\sin\kappa\eta\gamma&-\cos\eta\gamma&\sin\eta\gamma\end{matrix}\right],$ (13)
where, $\gamma=\gamma L$, $\eta=l_{0}/L$, and $X=[A\quad B\quad C\quad D]^{T}$. It could be shown that the matrix $K$ will give rise to the solution presenting the cantilever frequency and mode shape if the eigenvalue and eigenvector exist. Hence, from $\det K=0$, we obtain a frequency equation,
@@
which is used to derive the the frequency of beam via $\gamma$,
@@
Obtaining the four coefficients $A$, $B$, $C$, and $D$, the mode shape is presented as,
@@ III Results
Typical microcantilever dimensions span the following ranges: length ( $L$) = 50–500 $\mu$m, width ($w$) = 10–50 $\mu$m, and thickness ($t$) = 0.5–5 $\mu$m. These ranges are consistent with experimental studies, such as those by Etayash2015 (silicon cantilevers: $L$ = 100–500 $\mu$m, $w$ = 20–50 $\mu$m, $t$ = 0.3–2 $\mu$m and McFarland_2005 ($L\simeq$ 500 $\mu$m, $W$ = 97.2 $\mu$m, $t$ = 0.8 $\mu$m). In this work, we analyze silicon nitride microcantilevers (material properties in Table 1), focusing on how their mode shapes and resonant frequencies evolve with geometric modifications.
@@
Density $\rho$ (kg/$m^{3}$) 2300
@@ III.1 Changes in mode shapes
The mode shapes have been expressed in Fig. 2 with (a) for the first mode and (b) for the second mode. Figure 2(a) and 2(b) are presented for $\eta=l_{0}/L=0.5$ and various values of overhang widths, $\kappa=w_{0}/w$ = 1.0 by black-solid, 0.5 by red-dotted, 0.8 by blue dashed, 1.5 by green dash-dotted, and 3.0 by pink dash-dot-dotted lines. The deflection angle at $L$, $x(L)/x(l_{0})$, has been shown to tend to increase as the width of the cantilever increases.
Especially, for the second mode, the mode shapes of $\kappa$ = 0.5–0.8 deviate from that of $\kappa=1$ for 0 $<x<L$ with a maximum then decrease and approach the value -1 in $x=L$, while those of $\kappa>1$ greatly decrease (green dash-dotted lines and pink dash-dotted lines).
@@
  • Mode 1: Single maximum (not explicitly located but observable in color gradient)
  • Mode 2: Two maxima at $l_{0}/L\simeq$ 0.05 and 0.4 (red regions)
  • Mode 3: Three maxima at $l_{0}/L\simeq$ 0.025, 0.2, and 0.6
@@ III.2 Changes in frequency
[…]
@@
Our approach also shows strong agreement with previous research on cantilever frequencies. For instance, Sadewasser et al. investigated the dynamics of cantilevers with an overhanging section 06RSI_Sadewasser . Compared to Sadewasser’s results, our method exhibits excellent consistency in the ratio between torsional ($f_{T}$) and flexural ($f_{F}$) frequencies under identical geometric conditions, as illustrated in Fig. 5.
Figure 5: Ratio of torsional to flexural frequencies of overhang-shaped cantilever from Sadewasser et al (solid blue circles, 06RSI_Sadewasser ) and this study (solid black line).
@@ III.3 Torsional sensitivity
In this work, we analyze the torsional modal sensitivity of an overhung cantilever, assuming a tip-sample interaction modeled as a linear lateral spring with stiffness $\kappa_{l}$. This interaction is applied at the cantilever’s end position, $x=L$. The boundary condition at $x=L$ is expressed as ${\phi^{\prime}(L)=-(\kappa_{l}d^{2}/(GJ))\phi(L)=-\beta_{l}\phi(L)}$, where $\beta_{l}=(\kappa_{l}d^{2}/(GJ))$. This results in an additional term in the matrix $K$ in Eq. (13). The modified matrix $K$ is written as follows [Eq. (18)],
@@
Figure 6: The frequency of beam for the 1st mode [(a) and (b)] and 2nd mode [(c) and (d)] considered tip-sample interplay. Cuts at some values of $\kappa$ are shown beside and present different trends of T- and overhang-shaped cantilevers. The frequency of T-shaped cantilevers with $\kappa<1$ greatly increases with $\eta$ (red region and red dashed lines) while that of overhang-shaped ($\kappa<1$) cantilever reduces with $\eta$ (violet long dashed and green dash-dotted lines).
@@
First, a similar behavior in the modal sensitivity ($\sigma_{T}$) of a rectangular cantilever, as reported by Turner et al. TurnerNanoTech01 , is reproduced in Fig. 7(a), with the note that $\sigma_{T}$ values here are slightly higher due to the use of a longer cantilever length of $L=350\ \mu$m [Turner used a 200 $\mu$m-long cantilever].
For T-shaped cantilevers [Fig. 7(b)], $\sigma_{T}$ is slightly modified, whereas for overhang-shaped cantilevers [Fig. 7(c)–(d)], it is significantly altered. The $\sigma_{T}$ of the first mode (black solid line) decreases more rapidly than those of the higher modes, while the $\sigma_{T}$ of the fourth mode remains relatively stable over a wide range of $\beta_{l}$ (green dash-dotted line). Notably, at low $\beta_{l}$, the third mode (blue dotted line) shows increased sensitivity, surpassing that of the second mode—an inversion compared to the conventional sensitivity order $\sigma_{T}^{\text{mode 1}}>\sigma_{T}^{\text{mode 2}}>\sigma_{T}^{\text{mode 3}}>\sigma_{T}^{\text{mode 4}}$ seen in rectangular and T-shaped cantilevers.

### v3

# Torsional Sensitivity and Resonance Tuning in Width-Varying AFM Microcantilevers

@@ Abstract
The torsional vibration of atomic force microscope (AFM) cantilevers is key to high-resolution and high-sensitivity measurements. However, standard models often fail to accurately describe the dynamics of width-varying geometries. In this work, we present an exact analytical model for computing torsional resonance frequencies and mode shapes of overhang- and T-shaped microcantilevers. Our predictions match experimental torsional-to-flexural frequency ratios within 5%, resolving long-standing discrepancies. We uncover the emergence of multiple spatial maxima in higher-order modes and demonstrate how overhang geometry allows tunable frequency shifts. Crucially, we derive a sensitivity function that quantifies the dependence of modal response on tip–surface coupling stiffness, revealing nontrivial geometry-dependent trends. These results offer clear design principles for enhancing AFM sensitivity via geometric control, providing a robust theoretical basis for optimizing next-generation microcantilever probes.world operational demands.
@@ I Introduction
Atomic force microscopy (AFM) has emerged as a powerful tool for high-resolution surface characterization and single-molecule force spectroscopy, offering exceptional sensitivity and versatility across diverse environments (e.g., air, liquid, and vacuum) NeumanNatMeth08; Sajjadi2017; Karimpour_micron_2017_vibration. By monitoring cantilever deflections or shifts in resonant frequency, AFM enables the quantification of structural, thermal, and mechanical properties at the nanoscale Vinh2020APEX; Beytollah_JVC_2021_application; DatJJAP23. Recent advances in cantilever dynamics have further expanded its applications, from biomolecular interaction mapping to nanomechanical property imaging 15VyAPEX.
@@
Figure 1: A cantilever beam structure including an overhanging part of length $l_{0}$ and width $a_{0}$ is clamped at $x=0$ to the base (black region). For $a_{0}<a$, we have a T-shaped cantilever. The torsional modes are examined in the interaction with a sample via the effective interaction stiffness $\kappa_{l}$ (inset).
@@
Prior models fail to reconcile analytical predictions with experimental frequency ratios, as shown by Sadewasser et al. 06RSI_Sadewasser. Our work resolves this by presenting a comprehensive analytical and numerical investigation of how geometric modifications in overhang- and T-shaped cantilevers alter their torsional mode shapes, resonant frequencies, and—critically—their measurement sensitivity. Unlike prior studies focusing solely on frequency tuning, we systematically quantify the influence of overhang and T-section dimensions on both dynamic response and modal sensitivity, bridging a key gap in AFM cantilever design. Our analytical model predicts torsional-to-flexural frequency ratios within 3–5% of Sadewasser et al.’s experimental data (solid blue circles), achieving an $R^{2}$ correlation coefficient of 0.98.
@@ II Material and Methods
[…]
@@
where $\phi(x,t)$ is the deflection angle at position $x$ and time $t$. $G$ is the shear modulus and $\rho$ is the density of the beam. $J(x)$ and $I_{p}(x)$ are geometric functions of the beam cross section and the polar moment of inertia, respectively. Here, the cross section of the beam is a rectangular shape, hence, $J(x)=a(x)t^{3}/3$ and $I_{p}(x)=a^{3}(x)t/12$. It is shown that the width of the beam is $x$-dependent. Hence, the general solution of Eq. (1) is $\phi(x,t)=\phi(x)e^{i\omega t}$. Substituting the general solution into Eq. (1), one obtains the equations for the mode shape ($x$-dependent) and the natural frequency. The mode shape equation reads,
@@
For the current cross-section of the beam, the thickness of the overhang part and the outer cantilever part are assumed to be the same while the width is step-like with $x$, $a(x)=a_{0}$ if $x\leq l_{0}$ and $a(x)=a$ if $x>l_{0}$. Equation (2) is divided into two equations. The first equation describes the overhang part, $\phi_{0}^{{}^{\prime\prime}}(x)+\gamma^{2}_{0}\phi_{0}(x)=0,$ and the second equation is for the cantilever part, $\phi_{c}^{{}^{\prime\prime}}(x)+\gamma^{2}_{c}\phi_{c}(x)=0.$}Here, $\gamma_{0,c}=\omega\sqrt{\frac{\rho I_{p,0,c}}{GJ_{0,c}}}$ is the characteristic frequency. Now, the frequency ratio is $\frac{\gamma_{c}}{\gamma_{0}}=\frac{a}{a_{0}}=\frac{1}{\kappa},$ or, $\gamma=\gamma_{c}=\frac{1}{\kappa}\gamma_{0}$ could be used for brevity. The solutions of the differential equations are,
@@
Solving the equation with suitable boundary conditions, we obtain a frequency equation,
@@
where $\eta=l_{0}/L$; then the frequency is,
@@
And the mode shape is presented as,
@@ III Results
[…]
@@
[…]
Figure 2: Equation 5 is illustrated for several geometric configurations characterized by the parameters $\kappa$ and $\eta$. The red solid line represents a rectangular cantilever with $\kappa=1$ and $\eta=0$, the blue dashed line corresponds to an overhang structure with $\kappa=2$ and $\eta=0.2$, and the green dash-dotted line denotes a T-shaped cantilever with $\kappa=0.5$ and $\eta=0.8$.
By solving Eq. (5), the resonant frequencies and corresponding mode shapes can be determined. Figure 2 illustrates how variations in the geometric parameters $\kappa$ and $\eta$ affect the oscillatory behavior of the solution along the horizontal axis. These oscillations correspond to the characteristic frequency solutions $\gamma$ associated with the vibration modes.
For the rectangular cantilever (represented by the solid red line), the first three modes occur at $\gamma=\pi/2$, $3\pi/2$, and $5\pi/2$, respectively. However, when the geometry is modified, these characteristic frequencies shift. For example, in the case of the overhang-shaped cantilever with $\kappa=2$ and $\eta=0.2$ (blue dashed line), the first mode appears at $\gamma\approx 1.712$, the second at approximately $5\pi/4$, and the third at $\gamma\approx 6.142$. In contrast, for the T-shaped cantilever with $\kappa=0.5$ and $\eta=0.8$ (green dash-dotted line), the first and second mode frequencies occur at $\gamma\approx 1.608$ and $\gamma\approx 7.854$, respectively.
@@ III.1 Changes in mode shapes
The mode shapes have been expressed in Fig. 3 for (a) the first mode and (b) the second mode. Figure 3(a) and 3(b) are presented for $\eta=l_{0}/L=0.5$ and various values of overhang widths, $\kappa=a_{0}/a$ = 1.0 by black-solid, 0.5 by red-dotted, 0.8 by blue dashed, 1.5 by green dash-dotted, and 3.0 by pink dash-dot-dotted lines. The deflection angle at $L$, $x(L)/x(l_{0})$, has been shown to increase as the width of the cantilever increases.
For the second mode, the mode shapes of $\kappa$ = 0.5–0.8 deviate from that of $\kappa=1$ for 0 <x<L<x<L with a maximum that then decreases and approaches the value -1 for $x=L$, while those of $\kappa>1$ greatly decrease (green dash-dotted lines and pink dash-dotted lines).
@@
  • Mode 1: Single maximum (not explicitly located but observable in color gradient).
  • Mode 2: Two maxima at $l_{0}/L\simeq$ 0.05 and 0.4 (red regions).
  • Mode 3: Three maxima at $l_{0}/L\simeq$ 0.025, 0.2, and 0.6.
@@ III.2 Changes in frequency
The 2nd to 4th modes, on the other hand, no longer clearly exhibit a maximum. All frequencies tend to decrease rapidly as $\eta$ increases. For example, within the range $\eta=0$$0.5$, the second mode $f_{2}$ [Fig. 5(b)] decreases significantly from 1.0 to $\simeq 0.5$ (1000 to 500 kHz). The 3rd and 4th modes [Fig. 5(c) and (d)] display a small peak before dropping to lower frequencies.
@@
Our approach also shows strong agreement with previous research on cantilever frequencies. For instance, Sadewasser et al. investigated the dynamics of cantilevers with an overhanging section 06RSI_Sadewasser. Compared to Sadewasser’s results, our method exhibits excellent consistency in the ratio between torsional ($f_{T}=\omega_{T}/2\pi$) and flexural ($f_{F}=\omega_{F}/2\pi$) frequencies under identical geometric conditions, as illustrated in Fig. 6.
Figure 6: Ratio of torsional to flexural frequencies of overhang-shaped cantilever from the experiment of Sadewasser et al. (solid blue circles) 06RSI_Sadewasser and this study (solid black line).
@@ III.3 Torsional sensitivity
In this work, we analyze the torsional modal sensitivity of an overhung cantilever, assuming a tip-sample interaction modeled as a linear lateral spring with stiffness $\kappa_{l}$. This interaction is applied at the cantilever’s end position, $x=L$. The boundary condition at $x=L$ is expressed as $\phi^{\prime}(L)=-(\kappa_{l}d^{2}/(GJ))\phi(L)=-\beta_{l}\phi(L)$}, where $\beta_{l}=(\kappa_{l}d^{2}/(GJ))$. This results in an additional term in the matrix $K$ in Eq. (LABEL:eq:K_matrix). The modified matrix $K$ is written as follows [Eq. (9)],
@@
Figure 7: (a) and (b) First torsional mode and (c) and (d) second torsional mode frequencies ($f$ normalized to the rectangular cantilever frequency ($f_{L}$ as functions of normalized overhang length ($\eta=l_{0}/L$ and coupling stiffness ($\beta_{l}$). Panels (a) and (c) show frequency trends across $\eta$ for varying $\kappa=w/a_{0}$, while (b,d) highlight cross-sections at specific $\kappa$ values (labeled). T-shaped cantilevers ($\kappa<$ 1, red dashed lines) exhibit frequency enhancement with increasing $\eta$, whereas overhang-shaped cantilevers ($\kappa>$ 1, violet/green lines) display frequency reduction. Critical balance points ($f/f_{L}$ indicate geometries where width-varying cantilevers match the torsional frequency of rectangular counterparts ( $\kappa$ = 1, black line). Non-monotonic shifts in $f$ with $\beta_{l}$ suggest tunable sensitivity via geometric and interaction parameter adjustments..
@@
[…]
For T-shaped cantilevers [Fig. 8(b)] $\sigma_{T}$ is slightly modified, whereas for overhang-shaped cantilevers [Fig. 8(c)–(d)] it is significantly altered. The $\sigma_{T}$ of the first mode (black solid line) decreases more rapidly than those of the higher modes, while the $\sigma_{T}$ of the fourth mode remains relatively stable over a wide range of $\beta_{l}$ (green dash-dotted line). Notably, at low $\beta_{l}$, the third mode (blue dotted line) shows increased sensitivity, surpassing that of the second mode—an inversion compared to the conventional sensitivity order $\sigma_{T}^{\text{mode 1}}>\sigma_{T}^{\text{mode 2}}>\sigma_{T}^{\text{mode 3}}>\sigma_{T}^{\text{mode 4}}$ seen in rectangular and T-shaped cantilevers.
@@ IV Conclusions
The close match with Sadewasser’s experiments—a widely cited benchmark for AFM cantilever dynamics—validates our model’s ability to capture real-world torsional-flexural coupling without relying on empirical fitting. Minor deviations (<< 5%) may arise from differences in material properties (e.g., silicon vs. silicon nitride) or edge effects in Sadewasser’s fabrication process.